# Exploration and Incentivizing Participation in Clinical Trials

### v6

@@ Abstract
Participation incentives a well-known issue inhibiting clinical trials. We frame this issue as a non-standard exploration-exploitation tradeoff: the trial would like to explore as uniformly as possible, whereas each patient prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the patients. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for clinical trials. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. Our results extend to heterogeneous agents.
@@ 1 Introduction
Clinical trials are a standard way to evaluate safety and efficiency of medical treatments. A paradigmatic design considers $n=2$ alternative treatments, a.k.a. arms, e.g., a new and an old drug for the same medical condition, or a new drug and a placebo. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patients. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if the outcomes are generated by an adversary.
@@
A well-known issue which inhibits clinical trials is the difficulty of recruiting patients. This issue may be particularly damaging for large-scale clinical trials for wide-spread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging. In more abstract terms, we have a non-standard variant of exploration-exploitation tradeoff: the clinical trial would like to explore as uniformly as possible, whereas each patient prefers the trial to exploit, and might not participate otherwise.222 In the standard variant of this tradeoff, an algorithm strives to converge on the best action, not randomize uniformly.
@@
Our objective is to incentivize participation while optimizing statistical performance. We leverage information asymmetry between the clinical trial and the patients. We use preliminary data from the ongoing trial to skew the random choice of arms in favor of a better treatment (i.e., towards exploitation), while retaining a sufficient amount of exploration. This causes the patients to prefer participation to their outside options which cannot use the trial’s data.
@@
We make two assumptions that are very common in economic theory: the mechanism (i.e., the clinical trial) has the power to (i) prevent undesirable information disclosure to the agents (i.e., the patients), and (ii) commit to following a particular policy. While these assumptions can be problematic in many application domains, they are quite plausible in clinical trials. Indeed, preventing any information leaks in an ongoing clinical trial is standard practice, to avoid biasing patients and/or doctors in favor of better-performing treatments (e.g., see Detry et al., (2012, pp. 26-30)). Moreover, the design of a clinical trial is typically required by regulation to be fixed in advance, spelling out all applicable “decision points”, and (essentially) it needs to be revealed to the patients to meet the standards of “informed consent” (see e.g., Arango et al., (2012) for background).
@@
While clinical trials with $n=2$ arms is the paradigmatic case, our results seamlessly extend to $n>2$. Clinical trials with $\geq 3$ arms are studied in biostatistics, e.g., (Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, e.g., (Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our modelling approach. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentives constraints. A clinical trial (mechanism) sequentially interacts with self-interested patients (agents) like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for clinical trials: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. (In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs.) If not for the incentives constraints, the mechanism would simply choose arms uniformly at random.
@@
The basic case posits that all agents are similar (homogeneous agents). The mechanism assigns a randomized arm to each agent, according to some policy that may depend on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (reflecting a common practice in clinical trials). Participation incentives are modeled as Bayesian Individual Rationality (BIR), a standard notion in economic theory. In the general case (heterogeneous agents), each agent is endowed with a type reflecting medical history and subjective preferences.
The mechanism needs to ensure participation (only) according to the agents’ beliefs. There is a curious mismatch between these beliefs and the adversarial environment allowed by our statistical model. The agents believe that the world is much simpler: stationary rather than adversarial, like in stochastic bandits. The agents’ Bayesian prior represents the pre-existing medical knowledge, and all differences in outcome distributions are explained by the agents’ public types. Such agents’ beliefs are fairly realistic and appear necessary for tractability.
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage creates information asymmetry which drives participation incentives, whereas the data from the main stage determines the counterfactual estimates.
@@
Our guarantees focus on the main stage, assuming that enough data is collected in the warm-up stage. The aforementioned data can also be collected either endogenously (i.e., by a mechanism in our model), or exogenously, e.g., using paid volunteers. The key point here is that a relatively small amount of data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. Endogenous implementation of the warm-up stage can rely on prior work on incentivized exploration (more on this in “Related work” and Lemma 4.4).
We consider three model variants: from homogeneous agents to heterogeneous agents to a mechanism with estimated type frequencies. While heterogeneous agents are much more challenging, treating homogeneous agents first allows us to present key concepts and techniques in a more lucid way. The extension with estimated type frequencies allows the mechanism to pay less attention to rare-but-difficult agent types. Many ideas carry over from one variant to the next.
@@
We obtain an optimal solution for each model variant. First, we put forward a benchmark: a hypothetical policy that observes the true “state of the world” (according to agents’ beliefs) and optimizes a certain objective. Then, we design a mechanism that attains a certain worst-case upper bound on the estimation error, stated in terms of this benchmark and scaling as $1/T$. Finally, we prove that no mechanism can achieve better estimation error, up to constant factors.
@@
Stationarity and adaptivity. All mechanisms we design are “stationary” throughout the main stage, using the same “per-round policy” in all rounds. This per-round policy maps the current agent’s type and the past observations from the warm-up stage to a distribution over arms. In particular, the policy ignores all data collected in the main stage; this is desirable, as synchronizing patients and adapting to data tends to be challenging in practice.333 For several reasons: treating the patients and/or observing the outcomes may take a long time, the doctors involved cannot easily coordinate with one another, and it just complicates the logistics. So, many/most patients may need to be treated in parallel. Accordingly, clinical trials typically do not adapt to data, or have just a few pre-determined “review points” when one can make a data-adaptive decision, with non-data-adaptive “stages” in between. Our mechanisms confine data-adaptivity to the warm-up stage, which is asymptotically very short compared to the main stage.
In contrast, our negative results apply to all mechanisms that are not data-adaptive over some time period $S$, but are unrestricted otherwise. The time period $S$ must comprise a constant (but possibly small) fraction of rounds. The per-round policies can change over time, can themselves be jointly chosen at random, and can depend on the past observations in all rounds $t\not\in S$. In particular, this includes the “multi-stage” mechanisms discussed in Footnote 3.
Discussion: significance. We initiate the study of incentivized participation in clinical trials. Our model connects a standard statistical objective with an information design framework. We obtain optimal results within our model, and accomplish that with relatively simple mechanisms. On a technical level, we connect the dynamic process and statistical estimators with a static information-design problem, zeroing in on the appropriate benchmarks for each problem variant. While our results are most meaningful for trials with a large number of participants, participation incentives would make such trials more common and realistic.
@@
Map of the paper. Section 2 spells out our model, focusing on the homogeneous agents. Section 3 provides preliminaries, particularly on per-round policies and statistical estimators. In Section 4, we work out the case of homogeneous agents, introducing key ideas with a more accessible exposition. Heterogeneous agents are treated in Sections 5, 6 and 7, starting with an overview; Section 7 considers an extension with estimated type frequencies. The technical machinery in Sections 4, 6 and 7 is introduced gradually, and heavily reused in the later sections. Finally, Section 7.2 investigates how the benchmark with estimated type frequencies mitigates rare-but-difficult agent types.
@@ 1.1 Related work
Incentivized exploration (IE) is a closely related model, motivated by recommendation systems. The platform strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations, and makes them incentive-compatible via information asymmetry. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
@@
The main difference concerns performance objectives. IE strives to maximize the total reward under time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because patients cannot observe or alter their treatments (but can choose whether to participate). [^4] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Mechanism’s power to control information and commit to a particular policy are substantial assumptions in IE (e.g., Bahar et al., (2016); Immorlica et al., (2020) are written to mitigate them), but they are very benign for clinical trials, as discussed above. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” to create information asymmetry (see Lemma 4.4).
@@
A variant of IE creates incentives via payments (e.g.,  Frazier et al.,, 2014; Han et al.,, 2015). However, this work also assumes time-invariant reward distributions and optimizes total reward. Moreover, it lets the agents observe full history, which is not the case for clinical trials.
@@
Economic design. Using information asymmetry to incentivize agents to take particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor: the designer chooses an information structure, rather than payoff-relevant actions directly, and designer’s utility is directly derived from the agents’ actions, unlike a complex statistical objective such as our estimation error.
@@
Statistics and machine learning. Absent economic incentives, our problem reduces to the design of clinical trials in bioinformatics and conterfactual estimation in machine learning. A detailed review of either literature is beyond our scope. Data-adaptive designs for clinical trials are not uncommon (e.g., see survey Chow and Chang,, 2008). However, while prior work used data-adaptivity to improve the trial’s statistical properties, we use it for participation incentives. Our mechanisms use a standard counterfactual estimator called inverse propensity score (IPS). While more advanced estimators are known (e.g.,  Dudík et al.,, 2014; Swaminathan and Joachims,, 2015; Wang et al.,, 2017), IPS is optimal in some worst-case sense (see Section 3), which suffices for our purposes.
@@
While our mechanism follows the protocol of multi-armed bandits (Slivkins,, 2019; Lattimore and Szepesvári,, 2020), it strives to choose arms uniformly at random, whereas a bandit algorithm strives to converge on the best arm. [^5] Under the standard objective of cumulative reward/loss optimization, as expressed by regret. More related is “pure exploration” (starting from Mannor and Tsitsiklis,, 2004; Even-Dar et al.,, 2006), where a bandit algorithm predicts the best arm under a stationary outcome distribution, after exploring “for free” (e.g., uniformly) for $T$ rounds. In contrast, our model allows adversarially chosen outcomes, and requires estimates for all arms and all outcomes. Note that bandit algorithms, whether regret-minimizing or “pure exploration”, are generally not geared to incentivize participation. Interestingly, both stationary and adversarially chosen outcome distributions are well-studied in bandits, as, resp., “stochastic” and “adversarial” bandits.

### v8

@@ Abstract
Participation incentives a well-known issue inhibiting randomized clinical trials (RCTs). We frame this issue as a non-standard exploration-exploitation tradeoff: an RCT would like to explore as uniformly as possible, whereas each patient prefers “exploitation”, i.e., treatments that seem best. We incentivize participation by leveraging information asymmetry between the trial and the patients. We measure statistical performance via worst-case estimation error under adversarially generated outcomes, a standard objective for RCTs. We obtain a near-optimal solution in terms of this objective: an incentive-compatible mechanism with a particular guarantee, and a nearly matching impossibility result for any incentive-compatible mechanism. We consider three model variants: homogeneous patients (of the same “type” comprising preferences and medical histories), heterogeneous agents, and an extension with estimated type frequencies.
@@ 1 Introduction
Randomized Clinical Trials (RCTs) are a standard way to evaluate safety and efficiency of medical treatments. [^1] We use “RCT” acronym in this sense throughout. Another common meaning, “randomized controlled trial”, is also consistent with our model, since one of the arms there is typically a “control”. A paradigmatic design considers $n=2$ alternative treatments, a.k.a. arms, e.g., a new and an old drug for the same medical condition, or a new drug and a placebo. Each patient $t$ is assigned an arm $a_{t}$ independently and uniformly at random. An outcome $\omega_{t}$ is observed and evaluated according to some predefined metric $f(\cdot)$. The goal is to use the observations to estimate the counterfactual averages $\tfrac{1}{T}\,\sum_{t\in[T]}f(\omega_{a,t})$, $a\in[n]$, where $T$ is the number of patients. Independent per-patient randomization over arms is essential to remove selection bias and ensure statistical validity. Uniform randomization is usually preferred, as it minimizes the estimation error. In particular, one can obtain strong provable guarantees even if each arm’s outcomes are generated by an adversary, rather than sampled from some fixed (arm-dependent) distribution. [^2] This corresponds to (oblivious) adversary in adversarial bandits (Auer et al.,, 2002).
@@
A well-known issue which inhibits RCTs is the difficulty of recruiting patients (e.g., see surveys Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). This issue may be particularly damaging for large-scale RCTs for wide-spread medical conditions with relatively inexpensive treatments. Then finding suitable patients and providing them with appropriate treatments would be fairly realistic, but incentivizing patients to participate in sufficient numbers may be challenging.
We focus on patients’ reluctance to participate that stems from their preferences among the available treatments (as opposed to other factors, e.g., logistical difficulties). We frame this issue as a non-standard variant of exploration-exploitation tradeoff: the RCT always prefers to explore as uniformly as possible, whereas each patient prefers to exploit (choose treatment that seems best), and might not participate otherwise.  [^3] The standard variant of this tradeoff has an algorithm balance exploration vs exploitation for its own sake, so as to converge on the best arm.
@@
We initiate the study of incentivized participation in RCTs. Our objective is to incentivize participation while optimizing statistical performance. We leverage information asymmetry between the RCT and the patients. We use preliminary data from the ongoing trial to skew the random choice of arms in favor of a better treatment (i.e., towards exploitation), while retaining a sufficient amount of near-uniform exploration. This causes the patients to prefer participation to any alternative treatment options they might have (which also don’t have access to the trial’s data).
@@
We make two assumptions that are common in economic theory: the RCT has the power to (i) prevent undesirable information disclosure to the patients, and (ii) commit to following a particular policy. While such assumptions can be problematic in other applications, they are realistic in RCTs. Indeed, it is a standard practice to not reveal the history of an ongoing RCT and the assigned treatment, to avoid biasing patients and doctors in favor of better-performing treatments (e.g., see Detry et al., (2012, pp. 26-30)). Moreover, the design of an RCT is typically required by regulation to be fixed in advance, spelling out all applicable “decision points”, and (essentially) it needs to be revealed to the patients to meet the standards of “informed consent” (Arango et al.,, 2012).
@@
While RCTs with $n=2$ arms is the paradigmatic case, our results seamlessly extend to $n>2$. RCTs with $\geq 3$ arms are studied in biostatistics, e.g., (Hellmich,, 2001; Freidlin et al.,, 2008), and are important in practice, e.g., (Parmar et al.,, 2014; Redig and Jänne,, 2015).
@@
Our modelling approach. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentives constraints. An RCT (mechanism) sequentially interacts with self-interested patients (agents) like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^4] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. If not for the incentives constraints, the mechanism would simply choose arms uniformly at random.
@@
Participation incentives are modeled as Bayesian Individual Rationality (BIR), a standard notion in economic theory. The mechanism commits to a some randomized policy that assigns an arm to each agent, depending on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (as discussed above). BIR states that each agent must prefer participation to her “outside option”, in expectation over her Bayesian beliefs.
Crucially, the agents believe in much simpler world than the adversarial one allowed by our statistical model. Essentially, they believe one’s outcome distribution is determined by the assigned arm and the medical history. Such beliefs are fairly realistic, representing pre-existing medical knowledge, and appear necessary for tractability.
We consider three model variants. To isolate the core issues of incentivized participation, we consider on the fundamental case of homogeneous agents with the same preferences and medical histories. The general case of heterogeneous agents endows each agent with a type: a (public) medical history which determines the outcome distributions and (private) preferences over these outcomes. The third variant provides the mechanism with estimated type frequencies; this allows to mitigate the harmful influence of rare-but-difficult agent types.
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage creates information asymmetry which drives participation incentives, whereas the data from the main stage determines the counterfactual estimates. [^5] In particular, the warm-up data is discarded for the counterfactual estimates. This is a part the price of incentivizing participation in a (much larger) main stage.
@@
Our guarantees focus on the main stage, assuming that enough data is collected in the warm-up stage. The aforementioned data can also be collected either endogenously (i.e., by a mechanism in our model), or exogenously, e.g., using paid volunteers. The key point is that a relatively small amount of data, determined by the prior and the utility structure, suffices to bootstrap the main stage for an arbitrarily large number of agents. Endogenous implementation of the warm-up stage can rely on prior work on incentivized exploration (more on this in “Related work” and Lemma 3.6).
@@
We obtain a near-optimal solution for each model variants described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward a benchmark: a hypothetical policy that observes the true “state of the world” according to agents’ beliefs [^6] A mapping from arms (or type-arm pairs) to outcome distributions. and optimizes a certain objective. Then, we design a mechanism that attains a worst-case upper bound on the estimation error, stated in terms of this benchmark and scaling as $1/T$. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. Conceptually, we connect the dynamic process and statistical estimators with a static information-design problem. Many ideas carry over from one variant to the next.
@@
Our mechanisms are simple in that their “sampling distributions” (the distributions from which the arms are sampled) do not adapt to the data collected in the main stage. This is desirable in practice.777 Because treating the patients and/or observing the outcomes may take a long time, and the doctors involved may find it difficult to coordinate with one another. Consequently, many/most patients may need to be treated in parallel. So, RCTs typically do not adapt to data, or have just a few pre-determined “review points” when one can make a data-adaptive decision, with non-data-adaptive “stages” in between. In contrast, our negative results apply to a wide class of mechanisms, including the “multi-stage” mechanisms from Footnote 7. In particular, the sampling distributions can change over time and in “almost all” rounds they can depend arbitrarily on all past observations.888“Almost all” rounds here means: all rounds outside of a contiguous time interval comprising a constant (but possibly very small) fraction of rounds.
Discussion. Our contribution is not the mechanisms per se, but the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. While our results are asymptotic, and hence most meaningful for RCTs with a large number of participants, participation incentives would make such trials more common and realistic. For concreteness, we also illustrate our results with simple numerical examples (Appendix A).
We emphasize that theoretical participation incentives come at a price: the warm-up samples are discarded for counterfactual estimates, and the main-stage sampling distributions are skewed. Whether this price is acceptable in practice depends on how much these incentives actually help in recruiting patients. Ultimately, the “price of incentives” should be included in the “price of experimentation” that the society pays for the sake of obtaining reliable medical knowledge.
This paper is only concerned with patients’ initial decision to participate in an RCT. Their incentives to remain in the trial, e.g., if the treatment appears ineffective, is outside of our scope. [^9] This is indeed an issue in practice, as per (FDA,, 2001, Section 2.1.7.2). However, it is unclear if any approach based on information asymmetry could possibly help here: once the patient learns about the actual outcome, the “principal” does not have much/any informational advantage left.
@@
Map of the paper. Section 2 spells out our model, focusing on the homogeneous agents. Section 3 works out the case of homogeneous agents, introducing key ideas with a more accessible exposition. Heterogeneous agents are treated in Section 4; Section 5 considers an extension with estimated type frequencies. The technical machinery in Sections 3, 4 and 5 is introduced gradually, and heavily reused in the later sections.
@@ 1.1 Related work
Incentivized exploration (IE) is a closely related model, motivated by recommendation systems. The platform strives to incentivize users to explore when they would prefer to exploit. The platform issues recommendations (which the users can override), and makes them incentive-compatible via information asymmetry. Introduced in (Kremer et al.,, 2014; Che and Hörner,, 2018), IE is well-studied, see (Slivkins,, 2019, Chapter 11) for a survey. Most technically related are Mansour et al., (2020); Immorlica et al., (2019).
@@
The main difference concerns performance objectives. IE strives to maximize the total reward under time-invariant outcome distribution, or just explore each arm once. Neither serves our statistical objective, which requires independent near-uniform randomization in each round: reward-maximization pulls a mechanism in a very different direction, and merely exploring each arm $N$ times does not ensure statistical validity under adversarially chosen outcomes. The economic constraints are different, too, because patients cannot observe or alter their treatments (but can choose whether to participate). [^10] This distinction is consequential only for $n>2$ arms. Then the economic constraint in IE is stronger than ours. Mechanism’s power to control information and commit to a particular policy are substantial assumptions in IE (e.g., Bahar et al., (2016); Immorlica et al., (2020) are written to mitigate them), but they are very benign for RCTs, as discussed above. Hence, our mechanisms, analyses and technical guarantees are very different from those in IE. However, our mechanisms reuse IE for the “warm-up stage” (see Lemma 3.6).
@@
A variant of IE creates incentives via payments (e.g.,  Frazier et al.,, 2014; Han et al.,, 2015). However, this work also assumes time-invariant reward distributions and optimizes total reward. Moreover, it lets the agents observe full history, which is not the case for RCTs.
@@
Economic design. Using information asymmetry to incentivize agents to take particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor. The key distinction is that designer’s utility is (typically) directly derived from each agent’s actions. In contrast, our statistical objective depends on the sampling distributions and does not add up across agents.
Randomized Clinical Trials (RCTs). We only point out the key connections to the (huge) medical literature on RCTs, as a more detailed review is beyond our scope.
Difficulties in patient recruitment are widely acknowledged as a significant barrier (e.g.,  Adams et al.,, 1997; Grill and Karlawish,, 2010). Its reasons are well-studied, e.g., see surveys (Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). Patients preferences among treatments are listed among the major factors, along with other factors such as insufficient trust, logistical difficulties, or aversion to randomization and loss of control. Prior work suggested increasing the probability of the “new” treatment as a possible remedy (Vozdolska et al.,, 2009; Jenkins and Fallowfield,, 2000; Pocock,, 1979; Karlawish et al.,, 2008), albeit without a formal model of participation incentives and without theoretical guarantees. These papers focus on the scenario when the new treatment appears ex-ante preferable compared to non-participation, and do not revise the treatment probabilities as the new data arrives. Other suggested remedies, focusing on eligibility criteria, patient outreach, trial logistics, etc., are not as relevant to this paper.
Data-adaptive designs for RCTs are not uncommon, e.g., see Chow and Chang, (2008) for a research survey, and Detry et al., (2012); Pallmann et al., (2018) for practitioners’ guidance. However, while prior work used data-adaptivity to improve efficiency (the tradeoff between the trial’s size and statistical properties), we use it to create participation incentives. Accordingly, the novelty in our RCT design is in connecting it to the analysis of incentives (and the corresponding lower bounds on the statistical objective), not in the RCT design per se. As far as the design itself, our RCT for the homogeneous-agent case follows a very standard “epsilon-greedy” approach from multi-armed bandits, whereas the design for the heterogeneous agents appears new.
Our mechanisms are superficially similar to patient preference trials (PPTs) (Torgerson and Sibbald,, 1998), in that both allow patient preferences to impact treatment allocation. The key difference is that PPTs allow the patients to directly choose the chosen treatment (potentially at the expense of the internal validity), whereas in our setting a patient’s report only affects the sampling distribution (hence the internal validity is not compromised). Another distinction is that we elicit patient preferences over the treatment outcomes, not over the treatments themselves.
@@
Statistics and machine learning. Our mechanisms use a standard counterfactual estimator called inverse propensity score (IPS). While more advanced estimators are known (e.g.,  Dudík et al.,, 2014; Swaminathan and Joachims,, 2015; Wang et al.,, 2017), IPS is optimal in some worst-case sense (see Section 2.4 and Appendix D), which suffices for our purposes.
@@
While our mechanism follows the protocol of multi-armed bandits (Slivkins,, 2019; Lattimore and Szepesvári,, 2020), it strives to choose arms uniformly at random, whereas a bandit algorithm strives to converge on the best arm. [^11] Under the standard objective of cumulative reward/loss optimization, as expressed by regret. More related is “pure exploration” (starting from Mannor and Tsitsiklis,, 2004; Even-Dar et al.,, 2006), where a bandit algorithm predicts the best arm under a stationary outcome distribution, after exploring “for free” (e.g., uniformly) for $T$ rounds. In contrast, our model allows adversarially chosen outcomes, and requires estimates for all arms and all outcomes. Note that bandit algorithms, whether regret-minimizing or “pure exploration”, are generally not geared to incentivize participation. Interestingly, both “stationary” and adversarially chosen outcome distributions – like, resp., agent beliefs and statistical objective in our model – are well-studied in bandits, as, resp., “stochastic” and “adversarial” bandits.